### v1

@@ Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. A cut is relevant if and only if it is the a minimum weight cut between two distinct vertices.
@@ 1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanned the edges of $G$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [21]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [6, 13, 15] and structural biology [14]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [24], are of use [1].
@@ 3 Relevant Cuts
The following simple characterization of relevant elements holds for matroids in general [24]:
@@ 4 Computational Considerations
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. Even though there may be exponentially many relevant cuts, it is possible to encode them in an intermediary data structure that can be constructed in polynomial time. As shown in [9], the set all minimal $a,b$-cuts for all $a,b$, i.e., the set of relevant cycles, can be computed from a Gomory-Hu tree $T^{*}$ and the Picard-Queyranne (PQ) DAGs for the edges $st$ of $T^{*}$. The PQ DAG $D(s,t)$ is a compact representation of all minimal $s,t$-cuts [16], containing the pertinent cuts as the closed sets of the DAG. Both $T^{*}$ and the PQ-DAG $D(st)$ for any $st$ can be computed in polynomial time. From this combined data structure, $\mathcal{K}$ can be listed with $O(|\mathcal{K}|n)$ effort [9]. This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [24], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [24]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [18], and thus at least as hard as NP-complete decision problems [22].
@@
In addition to the approach of [9], several other algorithms for listing $\mathcal{K}$ are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract $\mathcal{K}$ using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [10]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [19] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
@@ Acknowledgements
This work was supported by the German Research Foundation (DFG) within SPP2363 (460865652) and the Novo Nordisk Foundation as part of MATOMIC (0066551). Research in the Stadler lab is supported by the German Federal Ministry of Education and Research BMBF through DAAD project 57616814 (SECAI, School of Embedded Composite AI).

### v2

@@ Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. We show that a cut is relevant if and only if it is the a minimum weight cut between two distinct vertices. Moreover, we give a characterization in terms of Picard-Queyranne DAGs that can be used to accelerate the enumeration of the relevant cuts.
@@ A Short Note on Relevant Cuts
Richard GolnikBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanythomas@bioinf.uni-leipzig.de
[…]
The Students of The Advanced Bioinformatics Course 2024Department of Computer Science, Leipzig University, Augustusplatz 10, D-04109 Leipzig, Germany
[…]
@@ 1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanned the edges of $G$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [24]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [8, 15, 18] and structural biology [17, 16]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [27], are of use [3].
@@ 3 Relevant Cuts
The following simple characerization of relevant elements holds for matroids in general [27]:
@@ 4 Algorithmic Considerations
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. Even though there may be exponentially many relevant cuts, it is possible to encode them in an intermediary data structure that can be constructed in polynomial time. As shown in [11], the set all minimal $a,b$-cuts for all $a,b$, i.e., the set of relevant cuts, can be computed from a Gomory-Hu tree $T^{*}$ and the directed acyclic graphs introduced by Picard and Queyranne [19] for the edges $st$ of $T^{*}$. A PQ-DAG $D_{s,t}$ is a compact representation of all minimal $s,t$-cuts . It is obtained from $G$ by contracting to a single node each connected component of the residual flow network for an arbitrary maximum $s,t$-flow. Each node in $D_{s,t}$ is identified with the subset of vertices in $V$ from which it was formed. The “closed” node sets in the DAG $D_{s,t}$, i.e., those without out-edges, correspond to the minimal $s,t$-cuts. Both the Gomory-Hu tree $T^{*}$ (for recent advances see [1, 2]) and the PQ-DAG $D_{s,t}$ for any pair $s,t$ can be computed in polynomial time. From this combined data structure, $\mathcal{K}$ can be listed with $O(|\mathcal{K}|n)$ effort [11].
This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [27], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [27]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [21], and thus at least as hard as NP-complete decision problems [25].
The enumeration of all relevant cuts can be sped up considerably with the help of observations made by Gusfield and Noar in [11]. First we need some more notation.
For any $u,v\in V$, we say that $v$ is a successor of $u$ in $D_{s,t}$, if either $u$ and $v$ belong to the same node in $D_{s,t}$ or if the node in $D_{s,t}$ that contains $v$ is a successor of the node in $D_{s,t}$ that contains $u$. Let $D_{s,t}(X,Y)$ with $X,Y\subseteq V$ and $X\cap Y=\emptyset$ denote the PQ-DAG resulting from $D_{s,t}$ by contracting the set $X$ with all its successors, and contracting the set $Y$ with all its predecessors.
Lemma 4.1.
Any closed set in $D_{s,t}(X,Y)$ that contains $s$ but not $t$ is a closed set in $D_{s,t}$ that contains $\{s\cup X\}$ but not $\{t\cup Y\}$.
Proof 4.2.
By construction, $s$ is contracted into a supernode with all its successors and $t$ is contracted into a super-node with all its predecessors to create $D_{s,t}$. Accordingly, the super-nodes containing $s$ and $t$ are the only vertices with out-degree 0 and in-degree 0, respectively. For any vertex $u\in V$, $s$ is a successor of $u$ and $t$ is a predecessors of $u$, as a consequence of this construction. Contracting $X$ with all its successors therefore includes $s$ but not $t$ as a successor. Analogously, contracting $Y$ with all its predecessors includes $t$ but not $s$ as a predecessors. Therefore, the super-node of out-degree 0 in $D_{s,t}(X,Y)$ represents $\{s\cup X\}$, and the super-node of in-degree 0 represents $\{t\cup Y\}$. It follows that any closed set in $D_{s,t}(X,Y)$ contains $\{s\cup X\}$ thereby $s$ but not $\{t\cup Y\}$ thereby $t$.
Lemma 4.3.
[11, Thm. 2.1] For a fixed pair of vertices $u,v\in V$, let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})$ be the $k$ minimum-weight tree edges along the path between $u$ and $v$ in $T^{*}$. A cut $(V,V^{\prime\prime})$ is a minimum $u,v$-cut if and only if it appears (as a closed set) in at least one PQ-DAG $D_{x_{i},y_{i}}(u,v)$ for some $1\leq i\leq k$, where $D_{x_{i},y_{i}}(u,v)$ is the graph resulting from contracting all successors of $u$ and all predecessors of $v$ in $D_{x_{i},y_{i}}$, respectively, into a single node.
Theorem 4.4.
A cut $D\in\mathfrak{D}$ is relevant if and only if it is represented by at least one closed vertex set in a PQ-DAG $D_{a,b}$ derived from an edge in the Gomory-Hu tree $T^{*}$, hence $\{a,b\}\in T^{*}$.
Proof 4.5.
$\Rightarrow$” By Thm. 3.6, if a cut is relevant, it is a weight-minimal $u,v$-cut for some pair of distinct vertices $u$ and $v$. By Lemma 4.3 any weight-minimal cut between two vertices $u$ and $v$ is represented by a closed set in exactly one PQ-DAG $D_{a,b}(u,v)$, where $ab$ is weight minimal along the path from $u$ to $v$ in $T^{*}$. By lemma 4.1 any $a$-$b$ cut in $D_{a,b}(u,v)$ is also a cut in $D_{a,b}$.
$\Leftarrow$” Each edge $e=\{a,b\}$ in a Gomory-Hu tree $T^{*}$ is weight minimal for at least the vertex pair $a$ and $b$ on the path consisting of only the edge $ab$ itself. The PQ-DAG $D_{a,b}$ represents all weight minimal cuts between $a$ and $b$ as closed sets. The closed sets are therefore weight minimal cuts for two distinct vertices $s$ and $t$ and hence relevant as a consequence of Thm. 3.6.
Therefore, it sufficient to compute only the PQ-DAGs associated to edges of the Gomory-Hu tree $T^{*}$. As the edges of a Gomory-Hu tree are labeled by the weights of associated minimal cuts, it is trivial to enumerate the relevant cuts in non-decreasing order using this method.
@@
In addition to the approach of [11] and its derivation, several other algorithms for listing $\mathcal{K}$ are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract $\mathcal{K}$ using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [12]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [22] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
@@
5 Computational Experiments
We implemented both methods derived from [11], enumerating closed sets in PQ-DAGs for all pairs of vertices (GUS-P) and only for PQ-DAGs for edges of $T^{*}$ (GUS-T), as well the algorithm for enumerating cuts by non-increasing weights proposed by [29] (YEH). The cut basis for $10,000$ individual chemical graphs from USPTO was computed for method on a consumer PC (Intel i9-9900K). Times were averaged over $5$ iterations and no parallelization was used. Graphs ranged in size between 4 and 198 nodes with a median of 30 at 3 to 211 edges and a median of 33. As such, chemical graphs are considered to be sparse. Unsurprisingly GUS-T always outperformed both GUS-P and YEH, see Table 1. Consistently, but more pronounced for larger graphs, YEH was significantly faster than GUS-P. Although GUS-P was the fastest algorithm for almost 1% of the graphs, manual inspection showed that this observation is most likely an artifact caused by fluctuations in loading times for very small graphs. Overall, YEH was a factor 2.2 slower then GUS-T to create a cut base. However, YEH is overall more versatile when enumerating cuts.
Table 1: Summary of benchmarked chemical graphs. GUS-T performs best for the vast majority of cases, with exception of few very small graphs.
Category GUS-T GUS-P YEH
Total Running Time [s] 8340.50 35035.97 18616.66
Fastest Method [# Graphs] 9916 84 0
@@ 6 Concluding Remarks
Availability
A reference implementation of the methods developed in this paper are available via GitHub https://github.com/TGatter/Cut-Base-Toolkit.
Author Contributions
ND, TG, RG, and PFS designed the study, obtained the theoretical results and drafted the manuscript. ND and TG created the reference implementation and executed the benchmarks. The students of the Advanced Bioinformatics Course 2024 implemented [29] as a part of a course project.
@@ Acknowledgements
This work was supported by the German Research Foundation (DFG) within SPP2363 (460865652) and the Novo Nordisk Foundation as part of MATOMIC (0066551). Research in the Stadler lab is supported by the German Federal Ministry of Education and Research BMBF through DAAD project 57616814 (SECAI, School of Embedded Composite AI). We gratefully acknowledge the contributions of the students of the Advanced Bioinformatics Course 2024: Marie Freudenberg, Julius Guntrum, Florian Kühn, Lukas Marche, Louisa Marie von Menges, and Kiet Ngo Tuan.